# Rebellious Student: A Complementary Learning Framework for Background Feature Enhancement in Hyperspectral Anomaly Detection

### v2

@@ Abstract
A recent class of hyperspectral anomaly detection methods that can be trained once on background datasets and then universally deployed—without per-scene retraining or parameter tuning—has demonstrated remarkable efficiency and robustness. Building upon this paradigm, we focus on the integration of spectral and spatial cues and introduce a novel “Rebellious Student” framework for complementary feature learning. Unlike conventional teacher-student paradigms driven by imitation, our method intentionally trains the spatial branch to diverge from the spectral teacher, thereby learning complementary spatial patterns that the teacher fails to capture. A two-stage learning strategy is adopted: (1) a spectral enhancement network is first trained via reverse distillation to obtain robust background spectral representations; and (2) a spatial network—the rebellious student—is subsequently optimized using decorrelation losses that enforce feature orthogonality while maintaining reconstruction fidelity to avoid irrelevant noise. Once trained, the framework enhances both spectral and spatial background features, enabling parameter-free and training-free anomaly detection when paired with conventional detectors. Experiments on the HAD100 benchmark show substantial improvements over several established baselines with modest computational overhead, confirming the effectiveness of the proposed complementary learning paradigm. Our code is publicly available at https://github.com/xjpp2016/FERS.
@@ I Introduction
To overcome these limitations and realize a parameter-free, training-free, and direct detection framework, Li et al. proposed the Anomaly Enhanced Transformation Network (AETNet) [14]. In AETNet, synthetic anomalies are injected into background HSIs to train a Transformer-based denoising autoencoder. Once trained, the model enhances anomalies in unseen HSIs without additional training, enabling conventional parameter-free detectors (e.g., RX) to achieve remarkable performance on large-scale benchmarks such as the HAD100 dataset [14]. However, its effectiveness heavily depends on data augmentation strategies, limiting its scalability.
To address this issue, we previously introduced Feature Enhancement via Reverse Distillation (FERD) [21]. Notably, FERD’s entire training process requires no data augmentation. Unlike AETNet’s anomaly enhancement approach, FERD focuses on background enhancement. Specifically, FERD is trained on background HSI datasets through a unified framework that integrates a reverse distillation mechanism [22] with a spectral feature alignment mechanism. The reverse distillation follows a student-teacher architecture [23, 24] that leverages an inverted information flow (comparable to an asymmetric encoder-decoder) to prevent overfitting, while the feature alignment strengthens background representation by enforcing spectral-domain consistency between the encoder and decoder features. By pruning redundant modules, FERD retains only a lightweight Feature Enhancement Network (FEN) that is strongly correlated with background spectral representation. Once trained, the FEN is capable of enhancing HSI background features, enabling parameter-free and training-free AD when combined with conventional detectors like RX. On the HAD100 dataset, FERD demonstrates superior efficiency and performance compared to AETNet. However, FERD’s design is primarily focused on spectral features, leaving spatial cues comparatively underutilized.
Figure 1: Schematic illustration of the proposed “rebellious student” framework. Unlike conventional student-teacher paradigms, our approach treats one branch as a rebellious student, which is encouraged to diverge from the teacher’s output and learn complementary information that the teacher overlooks.
@@
To overcome this limitation and fully exploit the complementarity between spectral and spatial information, we propose a fully decoupled stepwise training strategy that progressively enhances feature representations. This strategy unfolds in two successive stages: first enhancing spectral cues and then introducing spatial learning to capture information that the spectral branch alone cannot model.
To realize such “complementary learning,” we propose a non-traditional student-teacher framework. Unlike conventional “student-mimics-teacher” paradigms [23, 24], our design regards the spectral branch as the teacher and the spatial branch as a “rebellious student.” Instead of imitating the teacher, the student is explicitly encouraged to diverge, thereby learning complementary information. As illustrated in Fig . 1, this is achieved by driving the student’s output features to deviate from those of the teacher. Concretely, we enforce decorrelation between teacher and student features via cross-covariance and cosine similarity losses, while introducing variance-preserving regularization to prevent feature collapse. In addition, a reconstruction constraint fuses the two branches’ features to recover the original HSI, ensuring that the student learns meaningful spatial information rather than irrelevant noise.
@@
The overall training procedure is summarized in Fig. 2: Stage 1 — Spectral Branch Training: Leveraging FERD’s strong performance in enhancing background spectral features, the spectral branch inherits FERD’s training mechanism to obtain a spectral FEN specialized for background spectral representation enhancement. Stage 2 — Spatial Branch Training: In this stage, the spectral FEN is frozen and serves as the teacher network. The spatial FEN is trained as the rebellious student to learn background spatial features overlooked by the spectral FEN. Additionally, the spectral FEN is deliberately designed to compress spectral information, thereby explicitly suppressing spectral cues and promoting complementary spatial learning.
@@
    We propose a novel complementary learning paradigm, which extends FERD by explicitly integrating spectral and spatial information through a non-traditional student-teacher framework. In this design, the spatial branch is treated as a “rebellious student” that deliberately deviates from the teacher to capture complementary spatial cues.
@@
    We introduce a fully decoupled stepwise training strategy that first enhances spectral representations via FERD and subsequently trains a spatial branch with explicit suppression of spectral cues, ensuring progressive and complementary feature learning.
@@
    We provide comprehensive experimental validation and mechanistic insights: beyond demonstrating significant improvements over FERD on the HAD100 benchmark with minimal overhead, cross-scene experiments reveal that our enhancement reshapes feature distributions toward Gaussian-like forms with improved separability, explaining its universal benefits across both statistical and learning-based detectors
@@ III Proposed Method
This section introduces a novel “rebellious student-teacher” collaborative paradigm. The core idea is to fully exploit the complementarity between spectral and spatial features in HSIs, thereby enabling efficient background prior modeling and improving AD performance under parameter-free and training-free settings.
@@ III-B Stage 2: Training the Student Network (Spatial Branch)
    where $\text{CosSim}(\cdot,\cdot)$ is the cosine similarity function. The loss reaches its minimum value of 0 when the features are perfectly anti-correlated ($\text{CosSim}=-1$), and its maximum of 2 when they are perfectly aligned ($\text{CosSim}=+1$). By penalizing directional alignment, this loss ensures that the spatial features capture information in directions orthogonal to the spectral features.
@@ IV-B Experimental Settings
  • AETNet [14]: A parameter-free anomaly detection method, pioneeringly developed on HAD100, which enhances anomaly priors via data augmentation.

### v3

# Decoupled Complementary Spectral–Spatial Learning for Background Representation Enhancement in Hyperspectral Anomaly Detection

@@ Abstract
A recent class of hyperspectral anomaly detection methods can be trained once on background datasets and then deployed universally without per-scene retraining or parameter tuning, showing strong efficiency and robustness. Building upon this paradigm, we propose a decoupled complementary spectral–spatial learning framework for background representation enhancement. The framework follows a two-stage training strategy: (1) we first train a spectral enhancement network via reverse distillation to obtain robust background spectral representations; and (2) we then freeze the spectral branch as a teacher and train a spatial branch as a complementary student (the “rebellious student”) to capture spatial patterns overlooked by the teacher. Complementary learning is achieved through decorrelation objectives that reduce representational redundancy between the two branches, together with reconstruction regularization to prevent the student from learning irrelevant noise. After training, the framework jointly enhances background representations from both spectral and spatial perspectives, and the resulting enhanced features can be plugged into parameter-free, training-free detectors (e.g., the Reed–Xiaoli (RX) detector) for test-time deployment without per-scene retraining or parameter tuning. Experiments on the HAD100 benchmark demonstrate substantial improvements over representative baselines with modest computational overhead, validating the effectiveness of the proposed complementary learning paradigm. Our code is publicly available at https://github.com/xjpp2016/FERS.
@@ I Introduction
To overcome these limitations and enable a train-once, deploy-anywhere HAD framework—requiring no test-time retraining or per-scene parameter tuning—Li et al. proposed the Anomaly Enhanced Transformation Network (AETNet) [14]. AETNet injects synthetic anomalies into background HSIs to train a Transformer-based denoising autoencoder. Once trained, it enhances anomalies in unseen HSIs, allowing simple parameter-free detectors (e.g., RX) to achieve strong performance on large-scale benchmarks such as HAD100 [14]. However, its effectiveness depends heavily on data augmentation strategies, which may limit scalability.
To address this issue, we previously introduced Feature Enhancement via Reverse Distillation (FERD) [21], whose training requires no data augmentation. Unlike AETNet, which enhances anomaly priors, FERD enhances background representations by combining reverse distillation [22] with a spectral feature alignment mechanism, yielding a lightweight Feature Enhancement Network (FEN) that improves both efficiency and accuracy over AETNet on HAD100. Nevertheless, FERD is primarily tailored to spectral representations, leaving spatial cues underutilized.
Figure 1: Schematic illustration of the proposed decoupled complementary teacher–student framework. The spectral branch serves as a fixed teacher, while the spatial branch is trained as a complementary student (dubbed the “rebellious student”) to learn non-redundant cues beyond the teacher, guided by feature decorrelation objectives.
@@
To overcome this limitation, we propose a fully decoupled stepwise training strategy for complementary spectral–spatial background representation enhancement. The strategy proceeds in two stages: we first enhance spectral representations, and then introduce a spatial branch to capture cues that the spectral branch alone may miss.
Specifically, we adopt a teacher–student formulation where the spectral branch serves as a fixed teacher and the spatial branch is trained as a complementary student (the “rebellious student”). Rather than mimicking the teacher, the student is encouraged to learn non-redundant representations via feature decorrelation objectives (cross-covariance and cosine similarity penalties), together with variance-preserving regularization to avoid collapse. A reconstruction constraint further couples the two branches to recover the input HSI, ensuring that the learned spatial representations are meaningful rather than noise.
@@
The overall training procedure is summarized in Fig. 2: Stage 1 — Spectral Branch Training: we follow FERD to obtain a spectral FEN specialized for background spectral representation enhancement. Stage 2 — Spatial Branch Training: the spectral FEN is frozen as the teacher, and a spatial FEN is trained as the complementary student to learn spatial cues overlooked by the teacher. In addition, we deliberately compress spectral information in the student input to suppress spectral shortcuts and promote spatial learning.
@@
    We propose a novel complementary spectral–spatial learning paradigm that extends FERD by introducing a teacher–student formulation where a spatial branch is trained as a complementary student (dubbed the “rebellious student”) to learn non-redundant cues beyond the spectral teacher, guided by feature decorrelation objectives.
@@
    We introduce a fully decoupled stepwise training strategy that first learns a spectral enhancement network via FERD and then freezes it to supervise spatial learning with explicit suppression of spectral shortcuts, enabling progressive and complementary background representation enhancement.
@@
    We provide comprehensive experimental validation and mechanistic insights: in addition to consistent gains over FERD on HAD100 with modest overhead, cross-scene analyses show that the proposed enhancement regularizes background feature distributions toward more Gaussian-like structures and improves anomaly separability, helping explain its benefits for both statistical and learning-based detectors.
@@ III Proposed Method
This section presents a decoupled complementary spectral–spatial teacher–student paradigm for HSI background prior modeling, which boosts the performance of parameter-free, test-time training-free detectors (e.g., RX) without per-scene retraining or parameter tuning.
@@ III-B Stage 2: Training the Student Network (Spatial Branch)
    where $\text{CosSim}(\cdot,\cdot)$ is the cosine similarity function. The loss reaches its minimum value of 0 when the features are discouraging directional alignment (encouraging complementary directions) ($\text{CosSim}=-1$), and its maximum of 2 when they are perfectly aligned ($\text{CosSim}=+1$). By penalizing directional alignment, this loss ensures that the spatial features capture information in directions orthogonal to the spectral features.
@@ IV-B Experimental Settings
  • AETNet [14]: A test-time anomaly detection method, pioneeringly developed on HAD100, which enhances anomaly priors via data augmentation.